# Stimulating Imagination: Towards General-purpose Object Rearrangement

### v1

@@ Abstract
General-purpose object placement is a fundamental capability of an intelligent generalist robot, i.e., being capable of rearranging objects following human instructions even in novel environments. To achieve this, we break the rearrangement down into three parts, including object localization, goal imagination and robot control, and propose a framework named SPORT. SPORT leverages pre-trained large vision models for broad semantic reasoning about objects, and learns a diffusion-based 3D pose estimator to ensure physically-realistic results. Only object types (to be moved or reference) are communicated between these two parts, which brings two benefits. One is that we can fully leverage the powerful ability of open-set object localization and recognition since no specific fine-tuning is needed for robotic scenarios. Furthermore, the diffusion-based estimator only need to “imagine" the poses of the moving and reference objects after the placement, while no necessity for their semantic information. Thus the training burden is greatly reduced and no massive training is required. The training data for goal pose estimation is collected in simulation and annotated with GPT-4. A set of simulation and real-world experiments demonstrate the potential of our approach to accomplish general-purpose object rearrangement, placing various objects following precise instructions.
1 Introduction
@@
General-purpose object placement is a fundamental capability of an intelligent generalist robot. Much like humans, the robot must be capable of reasoning and recognizing target objects (even though it has never encountered before), and then constructing the rearrangement following human instructions. For example, if an instruction “put the spicy potato chips on the plate” is given, the ability of semantic understanding is required, i.e., reasoning about “spicy potato chips” even this phrase may be outside of the training distribution. Furthermore, the rearrangement should be physically-realistic by fully considering the physical structures, geometries and constraints.
@@
The great progresses of generative models provide researchers an insight of solving this challenging problem. Some of them introduce powerful models pre-trained on vision [3, 2, 7, 40], for initializing robotic policies or enhancing semantic understanding. Despite benefiting from the large-scale pre-training, it remains doubts about the generalization ability, since the amount of robotic manipulation fine-tuning data are far less than that encountered in a person’s experience (also less than pre-training). Another bottleneck is that these approaches do not look specifically at 3D spatial understanding. They assume that the underlying states of world can be characterized by images from certain angles.
@@
Some other works directly learn the object rearrangement skill in 3D space [29, 24, 23]. Specifically, the model takes point clouds as inputs and learns to directly estimate goal poses of rearranged objects. The training data are collected with physics simulators to ensure physically-valid results. However, given the fact that obtaining simulation data is expensive and time-consuming, the sizes of existing datasets are limited and scaling robot learning is difficult. Accordingly, these models lack the broad semantic understanding and reasoning ability for object localization, and may also fail to follow precise low-level instructions.
This paper presents an approach for leveraging pre-trained large vision models and 3D reasoning to enable general-purpose object rearrangement. The key insight is that the rearrangement process can be decoupled into three parts, namely object localization, goal imagination and robot control. We concentrate on the first two parts (the robot control can be done via Model Predictive Control or separated learned control policies [2]). Without loss of generality, we use “put the spicy potato chips on the plate” as an example. The moving object (spicy potato chips) and the reference object (plate) are first recognized and localized with a general-purpose image segmentation algorithm [20, 18], fully utilizing the powerful seasoning and semantic understanding ability derived from large-scale pre-training. The corresponding partial-view point clouds of these two objects are obtained with segments and RGB-D inputs. A natural language conditioned diffusion [14, 35] model is then introduced to “imagine” the goal poses that satisfy the rearrangement instruction and physical constraints. Note that the diffusion model can predict the object poses by only accessing their types, i.e., to be moved or reference, while without semantic information. This ease the burden of training, allowing us to learn well-generalized models with relatively little data.
@@
Figure 1: The proposed model can “imagine” the 3D poses of rearranged objects given a language instruction, being semantic-aware and physically-realistic.
In particular, the major contributions of this work are as follows.
@@
  • We demonstrate the potential for achieving general-purpose object rearrangement by decoupling the rearrangement process. The key is the ability of reasoning and recognizing various objects, as well as understanding 3D spatial relationships for a physically-realistic rearrangement.
@@
  • We present our approach, which leverages a powerful vision model pre-trained on large-scale data for broad object localization, and develops a diffusion-based model for physically-realistic 3D pose estimation. The results from both simulation and real-world experiments demonstrate the effectiveness of our approach, even with novel objects in novel environments.
@@
  • One key challenge is the lack of 3D object rearrangement data containing low-level instructions. We establish a GPT-assisted pipeline from a 3D perspective to generate high-quality data. The data is generated in simulation to ensure physical realism.
2 Related work
@@
3D scene generation. Vision generation models have endowed AI with visual imagination capabilities. Some research has attempted to apply diffusion models to arrangement tasks, thereby generalizing placement instructions. For instance, DALL-E-BOT [17] generates a goal image from text and matches it with an observed image to determine new object positions. However, it uses diffusion models directly, resulting in generated images that often differ from observed ones. DreamReal [16] employs a sampling strategy to sample candidate object positions in 3D space and uses Vision Language Models to score each candidate. Working in 3D space can make the generated scenes more physically realistic, avoiding collisions or placing objects in mid-air. StructFormer [24] and StructDiffusion [23] take a more direct approach, using transformers and diffusion models to edit observed 3D point clouds, enabling the execution of abstract instructions such as "set the dining table." However, they exhibit weak referential capabilities for objects.
Imitation Learning. Another research direction involves generating robotic actions directly. For example, Transporter [42] uses ResNet to generate robotic actions instead of object poses, while CLIP-PORT [34] combines CLIP [30] and Transporter to enhance object recognition capabilities. With the development of Multimodal Large Models, some works address robotic manipulation problems more end-to-end. Examples include Diffusion Policy [8], VIMA [15], and RT2 [3], which attempt to solve general robotic manipulation tasks by using language prompts and visual observations to generate visuo-motor actions directly. However, these imitation approaches usually require large amounts of robotic tele-operation or simulation data to achieve generalization capabilities.
3 Semantic-aware and physically-realistic object rearrangement
We introduce SPORT, a Semantic-aware and Physically-realistic Object RerrangemenT method. The pipeline of SPORT is illustrated in Figure 2.
Figure 2: The pipeline of SPORT. Given a rearrangement command, it takes RGB-D images of an initial scene as inputs, and generates the physically-realistic goal scene in 3D space.
3.1 Preliminaries and problem formulation
@@
Our key insight is that decoupling the components of rearrangement helps improve generalization. First, being capable of reasoning about novel objects and scenarios is necessary, even they may not be present in the robotic training data. A large language model (LLM) enhanced segment anything model (trained on Internet-scale data) is introduced to enable this. Specifically, the object types T =$=\{T_{1},T_{2},\ldots,T_{N}\}$ are inferred from W. Without loss of generality, we use subscripts $r$ and $m$ to represent the reference and moving objects, i.e., $O_{r}$ and $O_{m}$ respectively. The remaining objects, namely $\textbf{O}\setminus\{O_{r},O_{m}\}$, are irrelevant ones for the given command. The vision segmentation model takes $\textbf{I}_{rgb}$ and object descriptions as inputs and output segmentation masks of objects, denoted by $\textbf{M}=\{M_{1},M_{2},\ldots,M_{N}\}$. According to the M and $\textbf{I}_{d}$, the partial-view point clouds of objects can be derived, denoted by $\textbf{P}=\{P_{1},P_{2},\ldots,P_{N}\}$.
@@
Then the robot should act like a human that can “imagine” the 3D goal poses of objects after the rearrangement. A diffusion-based model is developed to achieve this, which takes P, W and T as inputs and estimate the pose $\mathbf{x}^{m}$ (only the object to be moved requires the pose estimation). One should note that the diffusion-based model only accesses to the object types (which object requires to be moved and which object is the reference) while without knowing what they are. “Moving $a$ to the left of $b$” and “moving $c$ to the left of $d$” are nearly equivalent to the pose estimation in our setting, because the final relative positions of objects are the same in these two commands. Though without semantic information, the physical validity of the rearrangement can still be ensured by understanding point cloud data. Benefiting from this, a well-generalized goal pose estimator can be trained without massive data.  [^1] After obtaining the goal poses, the robot control can be done via mature methods like MPC. Since this is not the focus of our work, related statements are not included.
3.2 SPORT for object rearrangement
We describe the details of the presented SPORT framework in this subsection.
@@ 3.2 SPORT for object rearrangement
Instruction parsing An LLM is utilized to understand and parse the natural-language command W. We still use the “put the spicy potato chips on the plate” as an instance. We prompt LLM, e.g., GPT [28, 5] or LLaMA [36, 37], to extract object types and corresponding descriptions: “spicy potato chips” to be moved and “plate” to be the reference, respectively.
@@
Object reasoning and segmentation An open-set segmentation model is then needed. We employ LISA [20] in this work though other similar models could also be used. LISA combines a multi-modal LLM (LLaVA [22]) with the segmentation decoder (SAM [18]), showing powerful capacity of complex semantic reasoning that requires world knowledge. For example, given an image containing two bags of potato chips, it can segment the spicy one according to the common knowledge “spicy snacks are usually packaged in red”.
@@
We do not fine-tune the vision model on robotic data. The reason is that the scale of robotic data is smaller than that of web data for pre-training large vision models [3, 40]. We can fully leverage the existing high-capacity of complex reasoning and semantic understanding, while fine-tuning may hurt the generalization. Furthermore, without the heavy work of fine-tuning, we can cost-effectively and flexibly use stronger models with the development of the community.
@@
Pose estimation We parameterize the 6-DoF pose as $(t,R)\in SE(3)$ (Special Euclidean Group). The goal pose estimator is based on a diffusion model, which is the basis of the recent remarkable AIGC approaches [31, 44]. It consists of several modules, i.e., a general-purpose text encoder, learnable type embeddings, a point cloud encoder and a vanilla transformer [38] as the backbone. We only use certain object (namely the moving one) to train the model, since the positions of the other objects remain unchanged after the rearrangement. The underlying idea behind this is similar to inpainting.
@@
Text encoder. We deploy BERT [10] as the text encoder along with its tokenizer, as it can understand general-purpose instructions in natural language form. Unlike previous works that may be limited to the tokens in their customized vocabulary [23], BERT is capable of broadly understanding various instructions and can well capture the information within the instructions. Though more powerful text encoder could be more helpful [32], we choose BERT as a trade-off between the need for strong semantic understanding and resource overhead.
@@
Type embedding. A set of learnable embeddings is introduced to indicate the token types, mainly the roles of corresponding objects in the rearrangement process, i.e., T. Four types are considered: the texts, the objects to be moved, the reference and irrelevant objects. Such embeddings help model to differentiate whether the poses of corresponding objects need to be changed.
@@
Object encoder. The object representations consist of two types of features. One encodes geometric and spatial information, and the other one encodes pose information at last time-step of diffusion model. For the former, we use a vanilla Point Cloud Transformer (PCT) model [13], given segmented partial-view point clouds of objects P. The mean position of the original point cloud is subtracted to ensure it does not retain any original pose information. For the latter, we use a multi-layer perceptron (MLP) to encode $(t,R)$. Apart from moving objects, the poses of other objects remain consistent with their initial pose. Finally, these two types of features are concatenated.
@@
Diffusion. A language-conditioned diffusion model is used to estimate the goal poses of objects. At each time-step, six types of embeddings (specifically the text, type, object, position, time and an extra token containing camera viewpoint information) are combined and fed to the backbone. The model then predict the poses at the current time-step, specifically the $t\in R^{3}$ and two vectors $a,b\in R^{3}$ to construct the rotation matrix $R\in SO(3)$. The position and time embeddings follow standard design, indicating the token positions in sequences and the time-step in diffusion, respectively. Padding is used to maintain a consistent number of input tokens. Note that the introduced extra token is essential, with which the model can effectively accomplish the instructions even when the camera viewpoint changes in real scenarios.
@@
Though all objects are included in model inputs, only the moving one gets involved in iterative pose estimations. It is because the diffusion model needs to know all object information to achieve relative positional movement and avoid collisions (thus all objects are required in the input), while only the pose of the moving object would change. Accordingly, we only add noise to the moving object pose during model training. The training objective can be formulated as
@@
where $\mathbf{\epsilon}$ is sampled from a standard normal distribution, $\mathbf{x}^{i}_{t}$ is the pose estimation of $i$-th object in O at $t$-timestep, and $\mathbb{I}_{i=m}$ is an indicator checking whether $i$-th object is the one to be moved.
3.3 GPT-assisted object rearrangement data generation

### v2

# Stimulating Imagination: Towards General-purpose “Something Something Placement”

@@ Abstract
General-purpose object placement is a fundamental capability of an intelligent generalist robot: being capable of rearranging objects following precise human instructions even in novel environments. This work is dedicated to achieving general-purpose object placement with “something something” instructions. Specifically, we break the entire process down into three parts, including object localization, goal imagination and robot control, and propose a method named SPORT. SPORT leverages a pre-trained large vision model for broad semantic reasoning about objects, and learns a diffusion-based pose es-timator to ensure physically-realistic results in 3D space. Only object types (movable or reference) are communicated between these two parts, which brings two benefits. One is that we can fully leverage the powerful ability of open-set object recognition and localization since no specific fine-tuning is needed for the robotic scenario. Moreover, the diffusion-based estimator only need to “imagine” the object poses after the placement, while no necessity for their semantic information. Thus the training burden is greatly reduced and no massive training is required. The training data for the goal pose estimation is collected in simulation and annotated by using GPT-4. Experimental results demonstrate the effectiveness of our approach. SPORT can not only generate promising 3D goal poses for unseen simulated objects, but also be seamlessly applied to real-world settings.
I INTRODUCTION
@@
General-purpose object placement is a fundamental capability of an intelligent generalist robot. This work focuses on the “something something placement” task: place some object relative to another one. Much like humans, the robot must be able to recognize target objects (even though it has never encountered before), and then rearranging the objects following human instructions. For example, if an instruction “put the spicy potato chips on the plate” is given, the ability of semantic understanding is required, i.e., reasoning about “spicy potato chips” even this phrase may be outside of the training distribution. Furthermore, the rearrangement should be physically-realistic by fully considering the physical structures, geometries and constraints.
@@
The great progresses of generative models provide an insight of solving this challenging problem. Some of them introduce powerful models pre-trained on vision [1, 2, 3, 4], for initializing robotic policies or enhancing semantic understanding. Despite benefiting from the large-scale pre-training, it remains doubts about the generalization ability, since the amount of robotic manipulation fine-tuning data are far less than that encountered in a person’s experience (also less than pre-training). Another bottleneck is that these approaches do not look specifically at 3D spatial understanding. They assume that the underlying states of world can be characterized by 2D images from certain angles.
@@
Some other works directly learn the object rearrangement skill in the 3D space [5, 6]. Specifically, the model takes point clouds as inputs and learns to directly estimate goal poses of rearranged objects. The training data are collected with physics simulators to ensure physically-valid results. However, given the fact that obtaining simulation data is expensive and time-consuming, the sizes of existing datasets are limited and scaling robot learning is difficult. Accordingly, these models lack the broad semantic understanding and reasoning ability for object localization, and may also fail to follow precise low-level instructions.
@@
Figure 1: SPORT can “imagine” the 3D poses of the rearranged objects given a language instruction, being semantic-aware and physically-realistic.
This paper presents a framework that leverages pre-trained large vision model and 3D spatial reasoning model to enable general-purpose object rearrangement with “something something” instructions. The key insight is that the process can be decoupled into three parts, namely object recognition and localization, goal imagination and robot control. We concentrate on the first two parts (robot control can be done via separated learned control policies [2] or Model Predictive Control). Without loss of generality, we use “put the spicy potato chips on the plate” as an example. The movable object (spicy potato chips) and the reference object (plate) are first recognized and localized by using an LLM-enhanced open-set segmentation algorithm [7, 8], fully utilizing the powerful seasoning and semantic understanding ability derived from large-scale pre-training. The corresponding partial-view point clouds of these objects are obtained based on segments and RGB-D inputs. A natural language conditioned diffusion [9] model is then introduced to “imagine” the goal poses in 3D space that satisfy the rearrangement instruction and physical constraints. Note that the diffusion model can predict the object poses by only accessing their types, i.e., to be moved or reference, while without semantic information. This ease the burden of training, allowing us to learn well-generalized models with relatively little data.
In particular, the contributions of this work are as follows.
@@
  • We demonstrate the potential to effectively address the “place something relative to something” task. The key is the ability of reasoning and recognizing various objects, as well as understanding 3D spatial relationships for a physically-realistic placement. Corresponding modules can be decoupled.
@@
  • We present our approach, which leverages a powerful vision model pre-trained on large-scale data for general-purpose object localization, and develops a diffusion-based model for physically-realistic pose estimation in 3D space. The results from both simulation and real-world experiments demonstrate the effectiveness of our approach, even on novel objects in novel environments.
@@
  • One key challenge is the lack of 3D object rearrangement data containing precise and low-level instructions. We establish a GPT-assisted pipeline from a 3D perspective to generate high-quality data. The data is generated in simulation to ensure physical realism.
II Related work
@@
3D scene generation. Vision generation models have endowed AI with visual imagination capabilities. Some research has attempted to apply diffusion models to arrangement tasks, thereby generalizing placement instructions. For instance, DALL-E-BOT [15] generates a goal image from text and matches it with an observed image to determine new object positions. However, it uses diffusion models directly, resulting in generated images that often differ from observed ones. Dream2Real [16] employs a sampling strategy to sample candidate object positions in 3D space and uses Vision Language Models to score each candidate. Working in 3D space can make the generated scenes more physically realistic, avoiding collisions or placing objects in mid-air. StructFormer [5] and StructDiffusion [6] take a more direct approach, using transformers and diffusion models to edit observed 3D point clouds, enabling the execution of abstract instructions such as ”set the dining table.” However, they exhibit weak referential capabilities for objects.
Imitation Learning. Another research direction involves generating robotic actions directly. For example, Transporter [17] uses ResNet to generate robotic actions instead of object poses, while CLIP-PORT [18] combines CLIP [19] and Transporter to enhance object recognition capabilities. With the development of Multimodal Large Models, some works address robotic manipulation problems more end-to-end. Examples include Diffusion Policy [20], VIMA [21], and RT2 [1], which attempt to solve general robotic manipulation with language prompts and visual observations to generate visuomotor actions directly. However, these imitation approaches usually require large amounts of robotic tele-operation or simulation data to achieve generalization capabilities.
Figure 2: The pipeline of SPORT. The inputs are the RGB-D images of the initial scene and the user command. The movable and reference objects are first extracted (“spicy potato chips” can be correctly obtained even with the distraction of chips in other flavors). Then a transformer-based diffusion model estimates the goal pose of the “spicy potato chips” in 3D space, satisfying the command and being physically-realistic.
III Semantic-aware and physically-realistic object rearrangement
We introduce SPORT, a Semantic-aware and Physically-realistic Object RerrangemenT method. The pipeline of SPORT is illustrated in Fig. 2.
III-A Preliminaries and Problem Formulation
@@
Our key insight is that decoupling the components of rearrangement helps improve generalization. First, being capable of reasoning about novel objects and scenarios is necessary, even they may not be present in the robotic training data. A large language model (LLM) enhanced segment anything model (trained on Internet-scale data) is introduced to enable this. Specifically, the object types T =$=\{T_{1},T_{2},\ldots,T_{N}\}$ are inferred from W. Without loss of generality, we use subscripts $r$ and $m$ to denote the reference and movable objects, i.e., $O_{r}$ and $O_{m}$ respectively. The remaining objects, namely $\textbf{O}\setminus\{O_{r},O_{m}\}$, are irrelevant ones for the given command. The vision segmentation model takes $\textbf{I}_{rgb}$ and object descriptions as inputs and output segmentation masks of objects, denoted by $\textbf{M}=\{M_{1},M_{2},\ldots,M_{N}\}$. According to the M and $\textbf{I}_{d}$, the partial-view point clouds of objects can be derived, denoted by $\textbf{P}=\{P_{1},P_{2},\ldots,P_{N}\}$.
@@
Then the robot should act like a human that can “imagine” the 3D goal poses of objects after the movement. A diffusion-based model is developed to achieve this, which takes P, W and T as inputs and estimate the 3D pose $\mathbf{x}^{m}$ (only the object to be moved requires the pose estimation). One should note that the diffusion-based model only accesses to the object types (which object requires to be moved and which one is the reference) while without knowing what they are. “Moving $a$ to the left of $b$” and “moving $c$ to the left of $d$” are nearly equivalent to the pose estimation in our setting, because the final relative positions of objects are the same in these two commands. Though without semantic information, the physical validity of the rearrangement can still be ensured by understanding point cloud data. Benefiting from this, a well-generalized goal pose estimator can be trained without massive data.  [^1] After obtaining the goal poses, the robot control can be done via mature methods like MPC. Since this is not the focus of our work, related statements are not included.
III-B SPORT for Something Something Placement
@@
Instruction parsing An LLM (e.g., GPT [22, 23] or LLaMA [24]) is utilized to understand and parse the natural-language command W. Still use “put the spicy potato chips on the plate” as an instance: “Spicy potato chips” is the object to be moved and “plate” is the reference.
@@
Object reasoning and segmentation An open-set segmentation model is then needed. We employ LISA [7] in this work though other similar models could also be used. LISA combines LLaVA [25] (multi-modal LLM ) with SAM [8] (segmentation decoder), showing powerful capacity of complex semantic reasoning that requires world knowledge. For example, given an image containing two bags of potato chips, it can segment the spicy one according to the common knowledge “spicy snacks are usually packaged in red”.
@@
We do not fine-tune the vision model on robotic data. The reason is that the scale of robotic data is smaller than that of web data for pre-training the large vision model [1, 4]. We can fully leverage the existing high-capacity of complex reasoning and semantic understanding, while fine-tuning may hurt the generalization.
@@
Pose estimation We parameterize the 6-DoF pose as $(s,R)$ $\in SE(3)$. The goal pose estimator is based on a diffusion model. Note that unlike recent AIGC approaches [26, 27], the diffusion model here is to generate the 6-DoF poses of the objects in the real-world, rather than the scene image directly. It consists of several modules, namely a general-purpose text encoder, learnable type embeddings, a point cloud encoder and a vanilla transformer [28] as the backbone. We only use certain object (the moving one) to train the model, since the other objects remain unchanged after the rearrangement. The underlying idea behind this is similar to inpainting.
@@
Text encoder. We deploy BERT [29] as the text encoder along with its tokenizer. Unlike several previous works that are limited to the tokens in their customized vocabulary [6], BERT can help the model to understand general-purpose instructions in natural language form.
@@
Type embedding. A set of learnable embeddings is introduced to indicate the token types in the placement, i.e., T. Four types are considered: the texts, the objects to be moved, the reference and irrelevant objects. With such embeddings, the model can differentiate whether the 6-DoF poses of corresponding objects need to be changed.
@@
Object encoder. There are two types of object representations (concatenated). One encodes geometric and spatial information by Point Cloud Transformer (PCT) [30], given segmented partial-view point clouds of objects P. The other one encodes pose information at the last time-step of diffusion model: using a multi-layer perceptron (MLP) to encode $(s,R)$. Apart from movable objects, the poses of other objects remain consistent with their initial pose.
@@
Diffusion. A language-conditioned diffusion model is used to estimate the goal poses of objects. At each time-step, six types of embeddings (specifically the text, type, object, position, time and an extra token containing camera viewpoint information) are combined and fed to the backbone. The model then predict the poses at the current time-step, specifically the $s\in R^{3}$ and two vectors $a,b\in R^{3}$ to construct the rotation matrix $R\in SO(3)$. The position and time embeddings follow standard design, indicating the token positions in sequences and the time-step in diffusion, respectively.
@@
All objects are included in model inputs, but only the one to be moved gets involved in iterative pose estimation. It is because the diffusion model needs to access all object information to achieve relative positional movement and avoid collisions, while only the pose of the movable object would change. Accordingly, we only add noise to the movable object pose during model training. Within the DDPM framework [9], the training objective can be formulated as
@@
where $\mathbf{\epsilon}$ is sampled from a standard normal distribution, $\mathbf{x}^{i}_{t}$ is the pose estimation of $i$-th object in O at $t$-timestep, and $\mathbb{I}_{i=m}$ is an indicator checking whether $i$-th object is the one to be moved. The generation process of goal pose is shown in Algorithm 1. It directly estimates the 6-DoF poses of the movable object in the 3D space.
Algorithm 1 3D Goal Pose Generation
1:$\mathbf{x}_{T}^{m}\sim N(\mathbf{0},\mathbf{I})$
[…]
3:     $\mathbf{z}\sim N(\mathbf{0},\mathbf{1})$ if $t$ > 1,  else $\mathbf{z}=\mathbf{0}$
4:     $\mathbf{x}_{t-1}^{m}=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}^{m}-\frac{1-\alpha_{t}}{\sqrt{1-\alpha_{t}}}\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t}^{m},t,\textbf{W},\textbf{T},\textbf{P}))+\sigma_{t}\mathbf{z}$
5:end for
6:return $\mathbf{x}_{0}^{m}$
III-C GPT-assisted Data Generation